From GlueXWiki
Jump to: navigation, search

May 7, 2015 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Gas puzzle
    • CO2 monitoring [2]
  2. CDC update (Beni, Mike)File:CDCEfficiencyUpdate7May.pdf
  3. FDC update (Lubomir)
  4. Electronics (Fernando, Cody, Chris, Nick)
  5. Engineering
  6. Other


Participants: Curtis, Naomi (CMU), Sean (NU), Luke, Fernando, Mike, Dave, Chris, Nick, Simon, and Lubomir (JLab).

Gas puzzle

- Last CO2 results linked above. CDC percentage was stable during the run at ~52-53% level. After the end of the run the sensor was moved to the tagger hall and connected to the spare package with 40% CO2, but it was showing 37-38%.

CDC update

- Mike showed results from the spring run for the CDC efficiencies (linked above) as function of distance to the wire and path length. Compared to the fall run and taken into account that the lower gain (due to the gas having more CO2) there's effectively increase in the efficiency thanks to the HV filter boxes. Two HV card showed some problems at the end of the run; Mike will look at this.

FDC update

- Lubomir compared the efficiencies from the fall and spring run, the latter with two HV settings 2200 and 2250, there's no visible change in the efficiency meaning no need to work at 2250V. Some of the bad channels in the third package started working by unknown reasons (last figure).


- Fernando proposed to start discussing if we can turn off some of the equipment during till the next run. CDC will take cosmics, then we will decide.